# organization=Université de Rennes, IRISA – UMR 6074,addressline=263 Av. Général Leclerc, city=Rennes, postcode=35000, country=France

### v1

@@ Abstract
In this article, we give an extended space formulation for the induced tree and path polytopes of chordal graphs with variables associated with the edge and vertex sets. Whereas the formulation for the induced tree polytope is easily seen to have a compact size, the system we provide for the induced path polytope has an exponential number of inequalities. We show which of these inequalities define facets and exhibit a superset of the facet-defining ones that can be enumerated in polynomial time. We show that for some graphs, the latter superset contains redundant inequalities. As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs .
Keywords— induced path , induced trees , extended formulations , polytope , chordal graphs
@@
All the graphs in this article are simple and connected unless specified. Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$. We respectively denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$. Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$. Given a subset of vertices $W\subseteq V$, let $E(W)$ denote the set of edges of $G$ having both endpoints in $W$ and $\delta(W)$ the set of all edges having exactly one endpoint in $W$. When $W$ is a singleton $\{w\}$, we will simply write $\delta(w)$. We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an extremity are said to be adjacent. For $F\subseteq E$, let $V(F)$ denote the set of vertices incident to any edge of $F$. Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$. When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$. Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$. Two vertices $u$ and $v$ are true twins if $N[u]=N[v]$ and false twins if $N(u)=N(v)$. A vertex $w$ is complete to a subset of vertices $W$ if it is adjacent to each vertex of $W$ and does not belong to $W$. A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal. A vertex is simplicial if its neighborhood induces a clique.
@@
An induced path (resp. hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp. $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ ). A subset of vertices induces a path (resp. hole) if its elements can be ordered into a sequence inducing a path (resp. hole). The length of a path or hole is its number of edges. A path or hole is called long if its length is at least 4. The parity of a path or hole is the parity of its length. Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior. A tree is a graph with no hole.
@@
The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$. For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
@@
A graph is $k$-chordal if its holes have maximum size $k$, 3-chordal graphs are called chordal. Equivalently, chordal graphs are those for which an order $(v_{1},\dots,v_{k})$ of its vertices exists such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$, such an order is called a perfect elimination order. Moreover, the maximal cliques of $G$ are among $N_{G_{i}}[v_{i}]$, which shows that they come in a linear number. A tree is a connected graph containing no holes.
@@
Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces. The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by dim $\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one. If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one. The polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is full dimensional when dim $\mathcal{P}$ is equal to $n$. In that case, a linear system $Ax\leq b$ that defines $\mathcal{P}$ is minimal if each inequality of the system defines a facet of $\mathcal{P}$. Moreover, these facet-defining inequalities are unique up to scalar multiplication. A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$. Equivalently, a face is the intersection of the supporting hyperplane of a valid inequality and $\mathcal{P}$. Extreme points of a polyhedron are its faces of dimension zero. Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality. A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is included in the $n$-dimensional hypercube and all its extreme points are binary. A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$.
@@
In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$. Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by the extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V}\times\{0,1\}^{E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0. Given a point = $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)_{G^{\prime}}$ its restriction to elements of $G^{\prime}$. A cone is a polyhedron described by a zero right-handed linear system; it is pointed if it contains no affine space. The elements of a cone can be described by conic combinations of finite sets of elements called generating. The minimal generating families of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators.
@@
In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral results associated with these problems.
@@
Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth]. Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made. A related polyhedra has, however, been characterized in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points are incidence vectors of paths between pairs of leaves in a tree: in trees, induced path and path are the same subgraphs, so the latter paper describes the induced path (between leaves) polytope of trees. Finally, extended space ILPs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
@@
In this paper, we first exhibit an extended formulation of compact size characterizing the induced tree polytope of chordal graphs, showing that the problem of computing a maximum vertex and/or edge weighted induced tree is solvable in polynomial time in chordal graphs. For a given graph, the induced tree polytope contains the induced path polytope, and both characterizations follow similar schemes of proof.
Even if the maximum vertex weighted induced path is already known to be solvable in polynomial time in chordal graphs [GAVRIL2002203, HANSEN2009135] it is unknown whether it is the case for vertex and edge weighted graphs. For this reason, we provide an extended formulation for the induced path polytope. The set of inequalities of this formulation is exponential, and we characterize which of them define facets. Finally, we show that a polynomial superset of the facet-defining ones can be enumerated efficiently. This exhibits a compact formulation for the induced path polytope of chordal graphs. Note that both latter results are necessary as our characterization of the facet-defining ones does not show that they come in a polynomial number.
@@ 1 Induced tree polytope of chordal graphs
In this section, we give an extended formulation for the induced tree polytope of a chordal graph. Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{T}(G)$ the following system of inequalities:
$\sum_{u\in K}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
@@
$x(V)-y(E)$ $=1$ (4)
We now prove that this system fully characterizes induced tree polytopes in chordal graphs.
@@
A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
Proof. $\Rightarrow"$ Note that the system of inequalities (1)–(3) defines a pointed cone $\mathcal{C}(G)$. The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}(G)$’s generators that satisfy (4). Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}(G)$ and satisfy (4). We show that they are the generators of $\mathcal{C}(G)$ by induction on a simplicial decomposition of $G$.
To initialize the induction, consider the graph $G$ with a single vertex, $\mathcal{C}(G)$ is the positive half line trivially generated by the unit vector.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that $\mathcal{C}(H)$ is generated by induced trees’ extended incidence vectors for $H=G\setminus\{v\}$. Let $(x^{*},y^{*})$ be a point of $\mathcal{C}(G)$, $(x^{*},y^{*})_{H}$ is a point of $\mathcal{C}(H)$ by construction. By the induction hypothesis, let $(x^{*},y^{*})_{H}$ be generated by the induced trees $\Gamma$ with coefficients $\lambda$. The proof consists of iteratively modifying $\Gamma$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Gamma$ and $\lambda$. Each operation preserves the value of $(x^{\prime},y^{\prime})_{H}$ but increments either $y_{uv}^{\prime}$ or $x_{v}^{\prime}$. We first modify the conic combination so that $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in E$. We consider the following notation $\lambda(\Gamma^{\prime})=\sum_{T\in\Gamma^{\prime}}\lambda_{T}$ for every $\Gamma^{\prime}\subseteq\Gamma$. Let us consider the following subsets of $\Gamma$:
@@
We refer to Figure 1 for an illustration of the sets of induced trees considered. Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$, we provide two different corresponding induced trees in different shades of green, which are the two bottom ones.
To each element $T$ of $\Gamma^{u}$, $T\cup\{v\}$ is an induced tree of $G$. This makes it possible to apply the Algorithm (1) over each element of $\bigcup_{u\in N(v)}\Gamma^{u}$, incrementing $y_{uv}^{\prime}$ and $x_{u}^{\prime}$.
Let $T$ be an element of $\Gamma^{u}$
$\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$
$T^{\prime}=T\cup\{v\}$
$\lambda_{T}=\lambda_{T}-\alpha$
$\lambda_{T^{\prime}}=\lambda_{T^{\prime}}+\alpha$
$T$ is removed from $\Gamma$ if $\alpha=\lambda_{T}$
$T^{\prime}$ is added to $\Gamma$
Algorithm 1 Reduction of $\Gamma^{u}$
Algorithm 1 increases $y_{uv}^{\prime}$ unless $y_{uv}^{\prime}=y_{uv}^{*}$ or $\Gamma^{u}$ is empty.
Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}(G)$ we have the following:
@@
The first equality is given by construction, the inequality is given by inequality (1) associated with $u$, and the remaining equalities are obtained by construction. This show how $y_{uv}^{*}=y_{uv}^{\prime}$ can be achieved on all edges $uv\in\delta(v)$. Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ and exhibit a conic comination of $(x^{*},y^{*})$.
@@
The induced tree polytope of a graph $G$ in the extended variable space is equal to $\mathcal{T}(G)$ if and only if $G$ is chordal.
The set of inclusion-wise maximal cliques of chordal graphs is polynomial (1)–(3), implying that Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.

### v2

# Extended formulations for induced tree and path polytopes of chordal graphs

@@ Abstract
In this article, we give two extended space formulations, respectively, for the induced tree and path polytopes of chordal graphs with vertex and edge variables. These formulations are obtained by proving that the induced tree and path extended incidence vectors of chordal graphs form Hilbert basis. This also shows that both polytopes have the integer decomposition property in chordal graphs.
Whereas the formulation for the induced tree polytope is easily seen to have a compact size, the system we provide for the induced path polytope has an exponential number of inequalities. We show which of these inequalities define facets and exhibit a superset of the facet-defining ones that can be enumerated in polynomial time. We show that for some graphs, the latter superset contains redundant inequalities. As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs.
keywords:
[…]
@@
All the graphs in this article are simple and connected unless specified. Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$. We respectively denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$. Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$. Given a subset of vertices $W\subseteq V$, let $E(W)$ denote the set of edges of $G$ having both endpoints in $W$ and $\delta(W)$ the set of all edges having exactly one endpoint in $W$. When $W$ is a singleton $\{w\}$, we will simply write $\delta(w)$. We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an endpoint are said to be adjacent. Given two subsets of vertices $W_{1}$, $W_{2}$, we denote by $\delta(W_{1},W_{2})$ the set of edges having one endpoint in both $W_{1}$ and $W_{2}$. For $F\subseteq E$, let $V(F)$ denote the set of vertices incident to any edge of $F$. Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$. When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$. Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$. Two vertices $u$ and $v$ are true twins if $N[u]=N[v]$ and false twins if $N(u)=N(v)$. A vertex $w$ is complete to a subset of vertices $W$ if it is adjacent to each vertex of $W$ and does not belong to $W$. A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal. Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$; we simply write $C(W)$ when clear from context. A vertex is simplicial if its neighborhood induces a clique.
@@
An induced path (resp. hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp. $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ ). Note that sets of vertices inducing cliques with three vertices are considered holes. An induced tree is a set of vertices inducing an acyclic graph. A subset of vertices induces a path (resp. hole) if its elements can be ordered into a sequence inducing a path (resp. hole). The length of a path or hole is its number of edges. A path or hole is called long if its length is at least 4. The parity of a path or hole is the parity of its length. Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior. A tree is a graph with no hole.
@@
The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$; parallel edges and loops are removed. For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
@@
A graph is $k$-chordal if its holes have maximum size $k$; 3-chordal graphs are called chordal. Equivalently, chordal graphs are those for which an order $(v_{1},\dots,v_{k})$ of its vertices exists such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$, such an order is called a perfect elimination order. Moreover, the maximal cliques of $G$ are among $N_{G_{i}}[v_{i}]$, which shows that they come in a linear number.
@@
The size of a linear system is its number of rows. Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces. The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by $\textrm{dim}~\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one. If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one. The polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is full dimensional when dim $\mathcal{P}$ is equal to $n$. In that case, a linear system $Ax\leq b$ that defines $\mathcal{P}$ is minimal if each inequality of the system defines a facet of $\mathcal{P}$. Moreover, these facet-defining inequalities are unique up to scalar multiplication. A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$. Equivalently, a face is the intersection of the supporting hyperplane of a valid inequality and $\mathcal{P}$. Extreme points of a polyhedron are its faces of dimension zero. Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality. A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is bounded and all its extreme points are binary. A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$. A cone is a polyhedron described by a linear system with a right hand side equal to 0; it is pointed if it contains no affine space. The elements of a cone can be described by conic combinations of finite sets of elements called generators. The minimal generating sets of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators they are in one to one correspondence with its face of dimension 1. A polyhedron $\mathcal{P}$ is said to have the integer decomposition property (IDP) if every integer point in the dilation $k\mathcal{P}$ for $k\in\mathds{Z}_{+}$ is obtained as the sum of $k$ elements of $\mathcal{P}$. A set of points $X=\{\chi_{1},\dots,\chi_{n}\}$ forms a Hilbert basis if every integer point in the conic hull of $X$ is expressible as an integer nonnegative combination of elements of $X$. The Hilbert basis and integer decomposition property are closely related. If the extreme points of a polytope $\mathcal{P}$ form a Hilbert basis, then $\mathcal{P}$ has the IDP.
@@
In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$. Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by its extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V\cup E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0. Given a point $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)|_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
In this paper, given a graph $G$, we consider the following polyhedra:
  • its induced tree cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
  • its induced tree polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
  • its induced path cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
  • its induced path polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
@@
In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral literature associated with these problems.
@@
Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth]. Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made. A related polyhedra has, however, been characterized in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points are incidence vectors of paths between pairs of leaves in a tree: in trees, induced path and path are the same subgraphs, so the latter paper describes the induced path (between leaves) polytope of trees. Finally, extended space integer linear programs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
@@
In this paper, we study both the induced tree polytope and the induced path polytope of chordal graphs in two separate sections. It turns out that beside being included one in the other, these two polytopes have similar properties. Each section starts by respectively showing that the induced path and tree extended incidence vectors of chordal graphs form a Hilbert basis of some cones, which we explicitly describe by linear systems of inequalities. We show that intersecting these cones with the correct hyperplane yields the induced tree and path polytope of chordal graphs in extended variable space. While the system we provide for the induced tree polytope is easily seen to be compact in chordal graphs, the one we provide for the induced path polytope has an exponential number of inequalities. Therefore, we characterize the ones that define facets and show that a superset can be enumerated in polynomial time. This implies that there exists a compact system describing the induced path polytope of chordal graphs. Finally, these results imply that the problems of finding a maximum vertex and edge weight induced tree or path are polynomially solvable in chordal graphs.
@@ 1 Induced tree polytope of chordal graphs
In this section, we show that the induced tree’s extended incidence vectors of chordal graphs form a Hilbert basis. Then we use this fact to exhibit a linear system describing the induced tree polytope of chordal graphs. Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{C}_{T}(G)$ the cone defined by the following system of inequalities:
$\sum_{u\in K\setminus\{v\}}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
@@
Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}_{T}(G)$.
@@
The induced tree extended incidence vectors of a chordal graph form a Hilbert basis.
Proof. We show that induced trees extended incidence vectors of a chordal graph $G=(V,E)$ form a Hilbert basis of $\mathcal{C}_{T}(G)$. As $\mathcal{C}_{T}(G)$ is defined with rational coefficients, each one-dimensional face contains integer points. This implies that a Hilbert basis of $\mathcal{C}_{T}(G)$ generates $\mathcal{C}_{T}(G)$.
We proceed by induction. For initialization, consider the graph $G$ with a single vertex. $\mathcal{C}_{T}(G)$ is the positive half-line trivially generated by the unit vector.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that the integer points of $\mathcal{C}_{T}(G\setminus v)$ are obtained as positive integer combinations of induced trees’ extended incidence vectors. Let $(x^{*},y^{*})$ be an integer point of $\mathcal{C}_{T}(G)$, $(x^{*},y^{*})|_{G\setminus v}$ is an integer point of $\mathcal{C}_{T}(G\setminus v)$ by construction. By the induction hypothesis, let $(x^{*},y^{*})|_{G\setminus v}$ be generated by the induced trees $\Gamma$ with integer coefficients $\lambda$. The proof consists of iteratively modifying $\Gamma$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Gamma$ and $\lambda$. For fixed $uv\in\delta(v)$, each operation preserves the value of $(x^{\prime},y^{\prime})|_{H}$ for $H=G\setminus uv$, and preserves nonnegativity and integrality of $\lambda$. We define the slackness $\Delta=\sum_{uv\in\delta(v)}\Delta(u,v)$ for $\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}$. We consider the following notation $\lambda(\Gamma^{\prime})=\sum_{T\in\Gamma^{\prime}}\lambda_{T}$ for every $\Gamma^{\prime}\subseteq\Gamma$.
Let us consider the following subsets of $\Gamma$, illustrated in Figure 1 :
@@
Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$, we provide two different corresponding induced trees in different shades of green, which are the two bottom ones.
We define the following operation to be applied on $(\Gamma,\lambda)$, which decreases $\Delta$ by an integer value.
$\mathbf{\Gamma^{u}}$-Reduction: let $T$ be an element of $\Gamma^{u}$. Set $\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$ increase $\lambda_{T\cup\{v\}}$ and decrease $\lambda_{T}$ by $\alpha$.
By construction, for any $T\in\Gamma^{u}$, $T\cup\{v\}$ induces a tree and that $\mathbf{\Gamma^{u}}$-Reduction preserves integrality of $\lambda$. Moreover, $\xi^{T}|_{H}=\xi^{T\cup\{v\}}|_{H}$ implying that $\lambda(\Psi)|_{H}$ is preserved.
$\mathbf{\Gamma^{u}}$-Reduction applies until $\Gamma^{u}$ is empty or $\Delta(u,v)=0$.
Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{T}(G)$ we have the following:
@@
The first equality is given by construction, the inequality is given by inequality (1) associated with $u$, and the remaining equalities are obtained by construction. This show how $\Delta(u,v)=0$ can be achieved on all edges $uv\in\delta(v)$. Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ with integer $\lambda$.
Note that Theorem 1.1 is not a characterization of the graphs for which the induced tree’s extended incidence vectors form a Hilbert basis. In fact, this statement is also true for $C_{4}$. However, it is straightforward to see that this property is closed by taking induced subgraphs. A natural question would be to characterize the graphs for which this property holds.
Theorem 1.1 implies the following.
Corollary 1.2.
The induced tree polytope and cone of a chordal graphs have IDP.
@@
Let us denote by $\mathcal{T}(G)$ the polytope defined as the intersection of $\mathcal{C}_{T}(G)$ and the following hyperplane:
$\{(x,y):x(V)-y(E)=1\}$ (4)
We now prove that $\mathcal{T}(G)$ characterizes induced tree polytopes in chordal graphs.
Theorem 1.3.
A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
Proof. $\Rightarrow"$ The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}_{T}(G)$’s generators that belong to (4). Theorem 1.1 concludes.
The set of inclusion-wise maximal cliques of chordal graphs is polynomial, implying that (1)–(3) has compact size. Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
@@
The induced tree polytope of a graph $G$ is equal to $\mathcal{T}(G)$ if and only if $G$ is chordal. Moreover, $\mathcal{T}(G)$ is described by a linear system of size $O(|V||\mathcal{K}|)$.